OpenStudy (anonymous):

can you please help me solve quadatric equation 4x^2-12x+9=0

OpenStudy (anonymous):

(3-2 x)^2 = 0 Factor x = 3/2

OpenStudy (anonymous):

\[4x^2-12x+9=(2x-3)^2=0\]\[(2x-3)^2 = 0\]\[(2x-3) = 0\]\[2x = 3\]\[x =3/2\]

OpenStudy (anonymous):

Never mind, I got it wrong. >___> Sorry. :'D

OpenStudy (amistre64):

factoring is great, when it can be done; but completing the square is the long version of the quadratic formula

OpenStudy (anonymous):

Why complete the square when you can factorise/use the formula?

OpenStudy (amistre64):

4x^2-12x+9=0 4(x^2 -3x+___) = -9 + 4(___) 4(x^2 -3x+9/4) = -9(4)/4 + 4(9/4) 4(x-3/2)^2 = 0 (x -3/2)^2 = 0 x = 3/2 maybe?

OpenStudy (amistre64):

completeing the square is just a good thing to know for other conics :)

OpenStudy (anonymous):

Exactly, no point doing extra work for nothing.

OpenStudy (amistre64):

and can be less of a headache when working the math

OpenStudy (amistre64):

unles syou do it by calculator :)

OpenStudy (anonymous):

Only if you are not sure about the formula, which completes the square for you